# Collapsing in polygonal dynamics

### v2

@@ 1 Introduction
A scaling symmetry is a rational group action $\rho:\mathbb{G}_{m}\times\mathcal{T}_{n}\to\mathcal{T}_{n}$ which commutes with the dynamic of $T$.
But some questions remain unanswered:
  1. 1.
    Which of the three eigenlines gives the collapse point, for what reason, and what is the meaning of the other ones ?
  2. 2.
    In [8, §9.2], Schwartz notices that for a family of transformations $\Delta_{k}$ (for which $\Delta_{1}$ is the pentagram map) and for certain polygons (called $k$-birds), there is a collapse behaviour like for the pentagram map with convex polygons. Moreover, the coordinates of collapsing are again given by a cousin of Glick’s formula. To quote him: “I wonder if this means that the collapse point exists for all starting points of the pentagram map. Even if the iterations go completely crazy under the map, perhaps they still collapse to the point predicted by Glick’s operator. The idea of a completely general collapse point has always seemed absurd to me, but maybe it is not. Nobody knows.
  3. 3.
    In [9, §7.2], Arnold and Arreche notice a similarity between the infinitesimal monodromy of the pentagram map and a so-called “hyperbolic barycenter” defined for the flat cross-ratio dynamics (a dynamical system for polygons on $\mathbb{P}^{1}(\mathbb{R})$, see [10, §6.2.1]). But the link is unclear.
In this paper, we answer these questions by considering the general setting of any polygonal dynamic with some additional hypotheses. From numerical experimentation, we conjecture that the collapsing generically [^4] By generic, we mean for a dense open subset of polygons, where the topology is given by the metric. occurs. This is reminiscent of an idea of Schwartz in his seminal paper about the pentagram map.
Conjecture 1.3 – 3.2.
Let $k$ be an algebraicly closed, complete field for the metric coming from a non-trivial valuation. Let $T$ be a polygonal transformation taking place in $\mathbb{P}^{d}(k)$, such that the dynamic on the moduli space is discretely integrable [^5] See definition 2.4..
Then for a generic choice of $n$-gon $P=((p_{1},\dots,p_{n}),M)$, there exists a matrix $\tilde{M}$ commuting with $M$, such that the vertices of $P$ collapse under the iteration of $T$ in the past/future towards $q_{-}$ and $q_{+}$, which are the repelling/attracting fixed points of $\tilde{M}$. Moreover, $q_{-}$ and $q_{+}$ depend continuously on $P$.
This brings an answer to question 1. We prove it in some specific configurations, for which we can exhibit some examples. It is important to remind it hold generically, and not for all polygons, since we find some non-generic counter-examples in subsection 6.2.
Theorem 1.4 – $3.$
Conjecture 3.2 holds for polygons whose dynamic is periodic on the moduli space.
Now, we want a formula for the collapse points. From lemma 3.1, we obtain that they have to be fixed points of the monodromy $M$ of the polygon $P$. However, we were originally interested in closed polygons only, and this is useless when $M=\mathrm{Id}$.
To resolve this issue, we use the scaling symmetry to define an infinitesimal monodromy $M^{\prime}$ (see lemma 2.1 which generalizes [7, Prop. 3.2]). This allows to retrieve the “true fixed points of the identity”, from which we obtain corollary 3.4.
Corollary 1.5 – 3.4.
Let $T$ be a polygonal transformation taking place in $\mathbb{P}^{d}(k)$ where $k$ is a complete non-trivially valued field, such that the dynamic is integrable on the moduli space and admits a scaling symmetry. Let $P$ be a generic closed polygon and $M^{\prime}$ be its infinitesimal monodromy.
Then the vertices of $P$ collapse under the iteration of $T$ in the past/future toward fixed point of $M^{\prime}$.
An other corollary is the generalization of [6, Thm 1.1]. Since the infinitesimal monodromy is a $(d+1)\times(d+1)$ matrix, finding its eigenlines only requires solving a linear system.
Corollary 1.6 – 3.5.
Let $P$ be a closed $n$-gon in $\mathbb{P}^{d}$ and $(p_{i}=[p_{i}^{(0)}:\dots:p_{i}^{(d)}])_{i=1,\dots,n}$ be its vertices. Let $q=[q^{(0)}:\dots:q^{(d)}]$ be a collapse point predicted by corollary 3.4. Then $q^{(0)},\dots,q^{(d)}$ are contained in a field extension of $\mathbb{Q}\big((p_{i}^{(j)})_{i=1,\dots,n}^{j=0,\dots,d}\big)$ of degree at most $d+1$.
This provides answers to questions 2 and 3 raised earlier.
Indeed, the pentagram map and its generalizations admit scaling symmetries and thus infinitesimal monodromies, so this answers Schwartz’s question about the reappearance of Glick’s formula.
This also explains the similarities observed by Arnold and Arreche between Glick’s formula for the pentagram map and the “hyperbolic barycenter” for cross-ratio dynamics. In fact, for any polygonal dynamic in $\mathbb{P}^{1}$ admitting a scaling symmetry, we give an explicit formula for the infinitesimal monodromy. Moreover, we give a degree $2$ polynomial equation that the collapse points have to verify, thus making an effective version of corollary 3.5.
Theorem 1.7 – 4.2.
@@
Suppose to have a polygonal dynamic on $\mathbb{P}^{1}$ admitting a scaling symmetry. Let $P$ be a closed $n$-gon, and $c_{i}(z)$ be the cross-ratio coordinates of $P_{z}$. Then the infinitesimal monodromy of $P$ is
$M^{\prime}=\begin{pmatrix}n/2+J&-K\\ I&n/2-J\end{pmatrix}$,
@@
$I=\sum_{i=1}^{n}\frac{c_{i}^{\prime}(1)}{c_{i}}\frac{1}{p_{i}-p_{i+1}},\quad J=\frac{1}{2}\sum_{i=1}^{n}\frac{c_{i}^{\prime}(1)}{c_{i}}\frac{p_{i}+p_{i+1}}{p_{i}-p_{i+1}},\quad K=\sum_{i=1}^{n}\frac{c_{i}^{\prime}(1)}{c_{i}}\frac{p_{i}p_{i+1}}{p_{i}-p_{i+1}}.$
$\chi_{P}(X,Y):=IX^{2}-2JXY+KY^{2}=0.$
This theorem also retrieves the invariant quantities $I,J,K,$ that appeared as particular cases in [10, §6.2] and [9]. Ideally, we could even retrieve an invariant pre-symplectic two form which keeps reappearing in each polygonal dynamics in $\mathbb{P}^{1}$, but we didn’t manage. The answer may come from a deeper interpretation in term of $\lambda$-lengths (see [10, §5.3]).
Finding the scaling symmetry of a system can be quite complicated. Once it’s done, we can apply theorem 4.2 to predict the collapse points. It is also the main tool to prove algebraic integrability (see [1] for an example on any algebraicly closed field). In this paper, we compute the scaling symmetry of the leapfrog map (defined in [11, §5.2]), the “flat” cross-ratio dynamic (studied in [10]), and the “staircase” cross-ratio dynamic. The last system originates from the “staircase” geometry in the theory of continuation of discrete holomorphic functions (see [12, §6] and references therein). However, it is more general than the initial problem since we allow to do the “flip” continuation on any vertices, regardless of any geometry. It makes up for a whole group of transformations, which we prove to be the affine symmetric group (see [13]). This group appeared in another polygonal dynamical system, the polygon recutting [3], which takes place in $\mathbb{R}^{2}$.
Outline of the paper: In section 2, we give some basic definitions and properties, and a review of known polygonal dynamics. In section 3 we formulate the conjecture 3.2 about generic collapsing. Then we focus on the closed polygons, for which corollary 3.4 predicts the collapse points. Section 4 is focused on polygonal dynamics in $\mathbb{P}^{1}$, for which theorem 4.2 gives an effective version of the corollary 3.4. We apply it in section 5 to some previously defined dynamics, and in section 6 to a new one, the staircase cross-ratio dynamics. We conclude by studying particular orbits of this system, which provides hints in favor of our conjecture.
2 Basic properties and examples
We are interested in polygonal dynamic $T$ taking place in $\mathbb{P}^{d}$ and admitting a scaling symmetry (see definitions 1.1 and 1.2). However, many reasonings would apply to more general cases (see [2, §2]). At the end of this section, table LABEL:tab:examples_polygonal_dynamics presents some known polygonal dynamics and their properties.
Let’s start by defining the main tool we’ll use for studying the dynamic on closed polygons.
Definition/Proposition 2.1.
Let $\rho$ be a scaling symmetry for a certain polygonal dynamic over $\mathbb{P}^{d}(k)$, where $k$ is a complete field.
Let $P]\in\mathcal{P}_{n}$ and $P]_{z}:=\rho_{z}\cdot[P]$. Up to a choice of a family of projective transformations, we can choose a smooth lift of the deformation
$P_{z}=((p_{1}(z),\dots,p_{n}(z)),M_{z})\in\tilde{\mathcal{T}}_{n}.$
Define the infinitesimal monodromy to be:
$M^{\prime}:=\left.\frac{dM_{z}}{dz}\right|_{z=1}.$
It is independent of the choice of lift and it is invariant under $T$.

### v3

@@ 2.1 Coordinate system in the moduli space
Let $P$ be an $n$-gon in $\mathbb{P}^{d}$. Its cross-ratio coordinates are defined to be $C=(c_{i})_{i\in\mathbb{Z}}$, where
$c_{i}=[c_{i,1}:\dots:c_{i,d+1}]:=C_{p_{i-1},p_{i},\dots,p_{i+d}}(p_{i+d+1}).$
They are $n$-periodic, and because of the non-degeneracy each $c_{i,j}$ is non-zero. We can normalize each $c_{i,d+1}$ to $1$, and the formula from proposition 2.4 gives us the values of the others $c_{i,j}\in\mathbb{P}^{1}\setminus\{0,\infty\}$.
Remark 3.
Other coordinate systems can be used to study polygonal dynamics, for instance the corner coordinates in the case of the pentagram map (see [ovsienkoPentagramMapDiscrete2010, p.415]). However, [Olver_2001, Thm. 3.10] tells us that any coordinate system is equivalent to the one given by cross-ratios. We will make concrete translation between different coordinates systems in section 5.
The cross-ratio coordinates $C$ parameterise the projective class $[P]$ of a non-degenerate polygon. With the following lemma, which is a generalization of [arnoldCrossratioDynamicsIdeal2022, Lemma 3.2], we can in fact reconstruct a polygon from the cross-ratio coordinates. Hence $\mathcal{T}_{n}$ is isomorphic to $(\mathbb{P}^{1}\setminus\{0,\infty\})^{dn}$. For the quotient space including degenerate polygons, see [weinreichGITStabilityLinear2024a].
Lemma 2.6.
Let $P$ be a $n$-gon in $\mathbb{P}^{d}$ and
$c_{i}=[c_{i,1}:\dots:c_{i,d+1}]$
be its cross-ratio coordinates. For any $i\in\{1,\dots,n\}$, define the matrix
$L_{i}=\begin{pmatrix}0&\dots&0&c_{i,1}\\ -c_{i,2}&\dots&0&c_{i,2}\\ \vdots&\ddots&\vdots&\vdots\\ 0&\dots&-c_{i,d+1}&c_{i,d+1}\end{pmatrix}\in\mathbb{P}\mathrm{GL}_{d+1}.$
Up to conjugation, suppose that $p_{0},p_{1},\dots,p_{d+1}$ is the standard frame. Then, the monodromy of $P$ is:
$M=L_{1}\dots L_{n}.$
Proof.
The proof follows as in $\mathbb{P}^{1}$. Take the lifts of $p_{0},p_{1},\dots,p_{d+1},p_{d+2}$:
$V_{0}=\begin{pmatrix}1\\ \vdots\\ 1\end{pmatrix},\quad V_{1}=\begin{pmatrix}1\\ 0\\ \vdots\\ 0\end{pmatrix},\quad\dots,\quad V_{d+1}=\begin{pmatrix}0\\ \vdots\\ 0\\ 1\end{pmatrix},\quad V_{d+2}=\begin{pmatrix}c_{1,1}\\ \vdots\\ c_{1,d+1}\\ \end{pmatrix}$.
A straightforward computation gives that
$L_{1}:\begin{cases}p_{0}\sim V_{0}&\mapsto c_{1,1}V_{1}\sim p_{1}\\ p_{j}\sim V_{j}&\mapsto-c_{1,j+1}V_{j+1}\sim p_{j+1}\quad\forall j\in\{1,\dots,d\}\\ p_{d+1}\sim V_{d+1}&\mapsto V_{d+2}\sim p_{d+2}\end{cases}.$
Moreover, because of proposition 2.3, we also have that
$L_{1}:p_{d+2}\sim V_{d+2}\mapsto V_{d+3}\sim p_{d+3}.$
Call $M_{i}=L_{1}\dots L_{i}$. By the exact same recurrence as in [arnoldCrossratioDynamicsIdeal2022, Lemma 3.2], we get that
$M_{i}(p_{0},\dots,p_{d+1},c_{i})\mapsto(p_{i},\dots,p_{i+d+1},p_{i+d+2}).$
Applying it to $M=M_{n}$ gives us the result. ∎
Computing the determinants of the $L_{i}$’s is easy, and yields the following corollary.
Corollary 2.7.
We have
$\mathrm{det}(M)=\prod_{i=1}^{n}\prod_{j=1}^{d+1}c_{i,j}.$
Since a polygonal transformation preserves the monodromy, the coefficients of the characteristic polynomial of $M$ are invariant through the dynamic. They are the usual candidates for Hamiltonians. We compute some of them in section 4.2.
2.2 Scaling symmetry and infinitesimal monodromy
Now that we have a description of the moduli space in term of cross-ratio coordinates $C:=(c_{i,j})_{1\leq j\leq d+1}^{1\leq i\leq n}$, understanding how a polygonal transformation $T$ acts revolves on finding a formula for $\tilde{C}:=T(C)$. Remember that $T$ is birational, so every $\tilde{c}_{k,l}$ is a rational expression in the $c_{i,j}$’s. [^7] With this observation, one could say that any polygonal dynamic is a cross-ratio dynamic. But this denomination is already devoted to specific systems.
Guessing the right equations linking $C$ with $\tilde{C}$ is quite tedious (see lemmas 5.4 and 6.5). When we get them, we can try to fit a so called spectral parameter. This leads to the following definition, more precise than 1.3. It can seem quite convoluted, see propositions 5.1, 5.5, 6.6 for examples.
Definition 2.8.
Let $T$ be a polygonal transformation. A scaling symmetry $\rho$ is a non-trivial action of the algebraic group $\mathbb{G}_{m}$ on $\mathcal{T}_{n}$,
$\rho:\begin{pmatrix}\mathbb{G}_{m}\times\mathcal{T}_{n}&\to&\mathcal{T}_{n}\\ (t,[P])&\mapsto&[P]_{t}\end{pmatrix}$,
which commutes with $T$.
Remark 4.
In the cross-ratio coordinates, $\rho$ expresses as a rational map in $C$ and $t$. Hence the action is not defined whenever one of the $c_{i,j}(t)$ is equal to 0 or $\infty$, which corresponds to a degenerate polygon.
Finding a scaling symmetry is a little miracle. There isn’t, for now, a method to produce them. It is the main tool to find Hamiltonians and prove integrability. It also produces a spectral curve, which is the base ingredient of algebraic integrability.
Remark 5.
A scaling symmetry acts on $\mathcal{T}_{n}$, but it can be lifted to $\tilde{\mathcal{T}}_{n}$. Indeed, if $P$ is an $n$-gon in $\mathbb{P}^{d}$, we deform its class to get $([P]_{t})$ and we lift it by defining
  • $P_{1}=P$,
  • $P_{t}$ verifies that $P_{t}]=[P]_{t}$, and that $(p_{0}(t),\dots,p_{d+1}(t))$ is constant.
By the first fundamental theorem of projective geometry, this lift exists and is unique. All the other lifts are given by $(A(t)\cdot P_{t})$, where $(A(t))$ is any algebraic family of projective transformations such that $A(1)=\mathrm{Id}$.
Remark 6.
A closed polygon $P$, when deformed through the scaling symmetry, gives a family of twisted polygons. This is the idea of “opening up” a polygon presented in [aboudLimitPointPentagram2022, Fig. 3]. Thanks to lemma 2.6, we can keep track of the monodromy $M(t)$, whose entries are rational in $C$ and $t$.
Since the coefficients of $M(t)$ are rational in $t$, we can formally differentiate it. This gives us our main tool.
Definition/Proposition 2.9.
Let $\rho$ be a scaling symmetry for a certain polygonal dynamic over $\mathbb{P}^{d}$.
Let $P$ be a closed polygon and $(P_{t})$ be a lift of the deformation by $\rho$. Let $M(t)\in\mathbb{P}\mathrm{GL}_{d+1}$ be the monodromy of $P_{t}$ and define the infinitesimal monodromy to be:
$M^{\prime}(1)\in\mathfrak{pgl}_{d+1}.$
It is independent of the choice of lift and invariant under the dynamic.
Remark 7.
Matrices in $\mathfrak{pgl}_{d+1}$ are defined up to affine transformation. Indeed, for any non-vanishing map $\lambda$, the families of matrices $(M(t))$ and $(\lambda(t)M(t))$ in $\mathrm{GL}_{d+1}$ represent the same projective transformations. Hence, we have:
$(\lambda M)^{\prime}(1)$ $=\lambda^{\prime}(1)\mathrm{Id}+\lambda(1)M^{\prime}(1).$
Proof.
It is exactly the same proof as in [aboudLimitPointPentagram2022, Prop 3.2], which treated the case of $\mathbb{P}^{2}(\mathbb{C)}$. Let $P_{t},\penalty 10000\ \tilde{P}_{t}$ be two lifts such that $P_{1}=\tilde{P}_{1}$, and $(A(t))$ be an algebraic family of projective transformations such that $\tilde{P}_{t}=A(t)\cdot P_{t}$. This forces $A(1)=\mathrm{Id}$ and $\tilde{M}(t)=A(t)M(t)A(t)^{-1}$. By differentiating, we get:
$\tilde{M}^{\prime}(1)=M^{\prime}(1)+\left[A^{\prime}(1),M(1)\right].$
Since $P_{1}$ is a closed polygon, we have $M(1)=\mathrm{Id}$ and so the commutator vanishes. Hence infinitesimal monodromy is well defined. [^8] As noted in [aboudLimitPointPentagram2022], the infinitesimal monodromy isn’t well defined for twisted polygons, since the only central element of $\mathbb{P}\mathrm{GL}_{d+1}$ is $\mathrm{Id}$.
The infinitesimal monodromy is defined from the family of monodromies $(M(t))$ coming from the scaling symmetry. But $T$ preserves the monodromy and commutes with the scaling symmetry, so it also preserves the infinitesimal monodromy. ∎
The infinitesimal monodromy can be explicitly computed. This is a generalization of Glick’s operator, first defined for the pentagram map [glickLimitPointPentagram2020] and then reinterpreted as an infinitesimal monodromy [aboudLimitPointPentagram2022].
Theorem 2.10.
@@ 2.2 Scaling symmetry and infinitesimal monodromy
Let $T$ be a polygonal transformation in $\mathbb{P}^{d}$ admitting a scaling symmetry. Let $P=(p_{1},\dots,p_{n})$ be a closed $n$-gon and $(c_{i,j})$ be its cross-ratio coordinates. Then, its infinitesimal monodromy is
$M^{\prime}(1):p\mapsto\sum_{i=1}^{n}\sum_{j=0}^{d}\frac{c_{i,j+1}^{\prime}(1)}{c_{i,j+1}(1)}\frac{\mathrm{det}(p_{i},\dots,p_{i+j-1},p,p_{i+j+1},\dots,p_{i+d})}{\mathrm{det}(p_{i},\dots,p_{i+d})}p_{i+j}.$
Proof.
We proceed as in the proof of [aboudLimitPointPentagram2022, Prop. 4.1], but with our cross-ratio coordinates. Let $P$ be a closed polygon. Up to a projective transformation, suppose that $(p_{0},\dots,p_{d})$ is the standard projective frame. Following remark 5, we deform $P$ into the family $(P_{t})$. For any $t\in\mathbb{G}_{m}$, we lift again the $p_{i}(t)$’s by choosing non-zero vectors $V_{i}(t)$’s as in the proof of lemma 2.6. Because of this lemma, we know that for any $t$ the monodromy is
$M(t)=L_{1}(t)\dots L_{n}(t).$
For any $i\in\{1,\dots,n\}$, define $M_{i}(t)=L_{1}(t)\dots L_{i}(t)$ and $M_{0}(t)=\mathrm{Id}$. Remark that
$L_{i+1}(t)\dots L_{n}(t)=M_{i}(t)^{-1}M(t).$
With this, we obtain that
$M^{\prime}(t)=(\sum_{i=1}^{n}M_{i-1}(t)L_{i}^{\prime}(t)M_{i}(t)^{-1})M(t).$
Since $M(1)=\mathrm{Id}$, this provides
$M^{\prime}(1)=\sum_{i=1}^{n}S_{i},\qquad\text{where }S_{i}:=M_{i-1}(1)L_{i}^{\prime}(1)M_{i}(1)^{-1}.$
Now we need to understand how each $S_{i}$ acts. Out of readability, we remove the “(1)(1)”. We have by the proof of lemma 2.6 that
$M_{i}:\begin{cases}V_{0}&\mapsto\alpha_{i,0}V_{i},\\ V_{j}&\mapsto\alpha_{i,j}V_{i+j}\quad\forall j\in\{1,\dots,d\},\\ V_{d+1}&\mapsto\alpha_{i,d+1}V_{i+d+2},\end{cases}$
@@
$\alpha_{i,j}=\begin{cases}1&\text{if }i=0\text{ or }j>d,\\ c_{i,1}\alpha_{i-1,1}&\text{if }j=0,\\ -c_{i,j+1}\alpha_{i-1,j+1}&\text{if }j\in\{1,\dots d\}.\end{cases}$ (1)
We also have
$L_{i}^{\prime}=\begin{pmatrix}0&\dots&0&c_{i,1}^{\prime}\\ -c_{i,2}^{\prime}&\dots&0&c_{i,2}^{\prime}\\ \vdots&\ddots&\vdots&\vdots\\ 0&\dots&-c_{i,d+1}^{\prime}&c_{i,d+1}^{\prime}\end{pmatrix}$
so it acts as
$L_{i}^{\prime}:\begin{cases}V_{0}&\mapsto c_{i,1}^{\prime}V_{1},\\ V_{j}&\mapsto-c_{i,j+1}^{\prime}V_{j+1}\quad\forall j\in\{1,\dots,d\},\\ V_{d+1}&\mapsto c_{i,1}^{\prime}V_{1}+\dots+c_{i,d+1}^{\prime}V_{d+1}.\end{cases}$
We assemble these statements to get that $S_{i}$ acts as:
$V_{i}$ $\stackrel{{\scriptstyle M_{i}^{-1}}}{{\mapsto}}$ $\frac{1}{\alpha_{i,0}}V_{0}$ $\stackrel{{\scriptstyle L_{i}^{\prime}}}{{\mapsto}}$ $\frac{c_{i,1}^{\prime}}{\alpha_{i,0}}V_{1}$ $\stackrel{{\scriptstyle M_{i-1}}}{{\mapsto}}$ $\frac{c_{i,1}^{\prime}\alpha_{i-1,1}}{\alpha_{i,0}}V_{i},$
$V_{i+j}$ $\mapsto$ $\frac{1}{\alpha_{i,j}}V_{j}$ $\mapsto$ $\frac{-c_{i,j+1}^{\prime}}{\alpha_{i,j}}V_{j+1}$ $\mapsto$ $\frac{-c_{i,j+1}^{\prime}\alpha_{i-1,j+1}}{\alpha_{i,j}}V_{i+j},$
where $j\in\{1,\dots,d\}$. Because of equation 1, we have that the coefficient in front of the output vectors are
$\frac{c_{i,j+1}^{\prime}}{c_{i,j+1}}\quad\text{for all }j\in\{0,\dots,d\}.$
Since the polygon $P$ is non degenerate, the family ($V_{i},\dots,V_{i+d}$) forms a basis of $\mathbb{A}^{d+1}$. Hence we can decompose any vector $V$ as
$V=\sum_{j=0}^{d}\frac{\mathrm{det}(V_{i},\dots,V_{i+j-1},V,V_{i+j+1},\dots,V_{i+d})}{\mathrm{det}(V_{i},\dots,V_{i+d})}V_{i+j},$
and by linearity we obtain
$S_{i}:V\mapsto\sum_{j=0}^{d}\frac{c_{i,j+1}^{\prime}}{c_{i,j+1}}\frac{\mathrm{det}(V_{i},\dots,V_{i+j-1},V,V_{i+j+1},\dots,V_{i+d})}{\mathrm{det}(V_{i},\dots,V_{i+d})}V_{i+j}.$
We projectivise it back and get that
$S_{i}:p\mapsto\sum_{j=0}^{d}\frac{c_{i,j+1}^{\prime}}{c_{i,j+1}}\frac{\mathrm{det}(p_{i},\dots,p_{i+j-1},p,p_{i+j+1},\dots,p_{i+d})}{\mathrm{det}(p_{i},\dots,p_{i+d})}p_{i+j}.$ (2)
Summing for all $i$’s gives the result. ∎
Remark 8.
As mentioned in remark 7, the infinitesimal monodromy is defined up to an affine transformation. This manifests in theorem 2.10 if we make a different choice of lift for the cross-ratio coordinates, or if we precompose the scaling symmetry $\rho$ with an automorphism of $\mathbb{G}_{m}$ (which are of the form $t^{k}$, with $k\in\mathbb{Z}^{*}$). The new coordinates are given by
$\tilde{c}_{i,j}(t)=\lambda_{i}(t)c_{i,j}(t^{k}),$
where the $\lambda_{i}$’s are non-vanishing functions. Hence we obtain:
$\frac{\tilde{c}_{i,j}^{\prime}(1)}{\tilde{c}_{i,j}(1)}$ $=\frac{\lambda_{i}^{\prime}(1)c_{i,j}(1)+\lambda_{i}(1)kc_{i,j}^{\prime}(1)}{\lambda_{i}(1)c_{i,j}(1)}$
$=\frac{\lambda_{i}^{\prime}(1)}{\lambda_{i}(1)}+k\frac{c_{i,j}^{\prime}(1)}{c_{i,j}(1)}.$
Plugging it in equation 2, we see that
$\tilde{S}_{i}=\frac{\lambda_{i}^{\prime}(1)}{\lambda_{i}(1)}\mathrm{Id}+kS_{i},$
so by summing over every $i$ we get
$\tilde{M}^{\prime}(1)=\left(\sum_{i=1}^{n}\frac{\lambda_{i}^{\prime}(1)}{\lambda_{i}(1)}\right)\mathrm{Id}+kM^{\prime}(1).$
2.3 Linear scaling symmetry
We restrain our study to a particular class of scaling symmetries.
Definition 2.11.
A scaling symmetry is said to be linear if the normalized cross-ratio coordinates are polynomials of degree at most one in $t$:
$c_{i,j}(t)$ $=c_{i,j}(1)+c_{i,j}^{\prime}(1)(t-1)$ $\forall(i,j)\in\{1,\dots,n\}\times\{1,\dots,d\},$
$c_{i,d+1}(t)$ $=1$ $\forall i\in\{1,\dots,n\}.$
So far, all known examples are linear. [^9] At first glance, this could seem to be wrong for the scaling symmetry of the pentagram map, defined in [ovsienkoPentagramMapDiscrete2010, Cor. 2.5] as $(x_{i},y_{i})\mapsto(tx_{i},t^{-1}y_{i}).$ However, the corresponding cross-ratio coordinates should be expressed in terms of $(x_{i},y_{i}^{-1})$, for which the scaling symmetry is linear. Maybe there is a reason for this. Following the spirit of [arnoldCrossratioDynamicsIdeal2022], we define for any $I\subset\{1,\dots,n\}\times\{1,\dots,d\}$ the quantities
$c_{I}:=\prod_{(i,j)\in I}c_{i,j}(1),\quad c^{\prime}_{I}:=\prod_{(i,j)\in I}c^{\prime}_{i,j}(1)$
A direct computation gives us the following lemma.
Lemma 2.12.
We have for any $I\subset\{1,\dots,n\}\times\{1,\dots,d\}$ and $t\in\mathbb{G}_{m}$:
$c_{I}(t)=c_{I}\sum_{J\subset I}\frac{c_{J}^{\prime}}{c_{J}}(t-1)^{|J|}.$
This gives an simple formula for the determinant.
Lemma 2.13.
Let $T$ be a polygonal transformation in $\mathbb{P}^{d}$ admitting a linear scaling symmetry. Let $P$ be an $n$-gon (twisted or closed) and $M(t)$ be the deformation of its monodromy from remark 5. For $0\leq k\leq nd$, define:
$R_{k}=\sum_{|I|=k}\frac{c_{I}^{\prime}}{c_{I}}.$
They are preserved by the dynamic and projective transformations, and verify:
$\mathrm{det}(M(t))=\mathrm{det}(M(1))\sum_{k=0}^{nd}R_{k}(t-1)^{k}.$